# UAV-Mounted IRS (UMI) in the Presence of Hovering Fluctuations: 3D Pattern Characterization and Performance Analysis

### v2

@@ Appendix A Proof of Lemma 1
$\mu_{\varepsilon_{\theta_{q}}}=A_{\theta_{qx}}\mu_{x}+A_{\theta_{qy}}\mu_{y},$ (92)
$\sigma^{2}_{\varepsilon_{\theta_{q}}}=A^{2}_{\theta_{qx}}\sigma^{2}_{x}+A^{2}_{\theta_{qy}}\sigma^{2}_{y}.$
@@ Appendix B Proof of Lemma 3
Given that, $Z_{x}$ and $Z_{y}$ are the RVs that are dependent on the fluctuations of UMI and follow a Gaussian distribution, the PDF of $g_{ax}$ will be given by
$f_{g_{ax}}=\sum_{i=0}^{lD_{a}-1}{p_{ax}(i)\times\delta(g_{ax}-\frac{D_{a}^{2}(1-cos(\frac{2\pi i}{D_{a}}))}{2\pi^{2}i^{2}})},$ (93)
@@
$p_{ax}(i)=Q(\frac{2i-D_{a}N_{q}\mu_{Z_{x}}}{D_{a}N_{q}\sigma_{Z_{x}}})$ $-Q(\frac{2(i+1)+D_{a}N_{q}\mu_{Z_{x}}}{D_{a}N_{q}\sigma_{Z_{x}}})$ (94)
$\qquad+Q(\frac{2i+D_{a}N_{q}\mu_{Z_{x}}}{D_{a}N_{q}\sigma_{Z_{x}}})$ $-Q(\frac{2(i+1)-D_{a}N_{q}\mu_{Z_{x}}}{D_{a}N_{q}\sigma_{Z_{x}}}).$
Note that, the PDF of RV $g_{ay}$ can be derived similarly to (93) by replacing the subscript $x$ with $y$. According to (14), $\mathbb{G}_{a}=g_{ax}\times g_{ay}$ and applying (93), the PDF of array factor, $G_{a}$ conditioned on the array factor in x-axis, $g_{ax}$ can be given as
@@
$\qquad\quad\sum_{i=0}^{lD_{a}-1}\frac{p_{ay}(i)}{g_{ax}}\times\delta(\frac{\mathbb{G}_{a}}{g_{ax}}-\frac{D_{a}^{2}(1-cos(\frac{2\pi i}{D_{a}}))}{2\pi^{2}i^{2}}),$
@@
[…]
$=\sum_{i=0}^{lD_{a}-1}\sum_{j=0}^{lD_{a}-1}{\int\frac{p_{ax}(i)p_{ay}(j)}{g_{ax}}}$
×$\quad\times\delta(\frac{\mathbb{G}_{a}}{g_{ax}}-\frac{D_{a}^{2}(1-\cos(\frac{2\pi i}{D_{a}}))}{2\pi^{2}i^{2}})$
×$\quad\times{\delta(g_{ax}-\frac{D_{a}^{2}(1-\cos(\frac{2\pi j}{D_{a}}))}{2\pi^{2}j^{2}})}dg_{ax}$
$=\sum_{i=0}^{lD_{a}-1}\sum_{j=0}^{lD_{a}-1}{p_{ax}(i)p_{ay}(j)\delta(\mathbb{G}_{a}-q_{ax}(i)q_{ay}(j)))},$
@@
$q_{ak}(i)=\frac{D_{a}^{2}(1-\cos(\frac{2\pi i}{D_{a}}))}{2\pi^{2}i^{2}},\quad k\in\{x,y\}.$ (97)
@@
$f_{\mathbb{G}_{a}}(\mathbb{G}_{a})=\sum_{i=1}^{l^{2}D_{a}^{2}}p_{a}(i)\delta(\mathbb{G}_{a}-q_{a}(i)),$ (98)
where $\mathbf{q}_{a}=\mathbf{q}_{ax}\otimes\mathbf{q}_{ay}$ and $\mathbf{p}_{a}=\mathbf{p}_{ax}\otimes\mathbf{p}_{ay}$. Here, $\mathbf{q}_{ak}\in\mathbb{R}^{lD_{a}\times 1}$ and $\mathbf{p}_{ak}\in\mathbb{R}^{lD_{a}\times 1}$ where $k\in\{x,y\}$. $p_{ak}(i)$ and $q_{ak}(i)$ are given in (94) and (97), respectively.
Appendix C Proof of Lemma 4
Since $\varepsilon_{\theta_{q}}$ is the random variable that relies on the fluctuations of IRS and follows a Gaussian distribution, the PDF of $g_{eq},~{}q\in\{t,r\}$ can be represented as
$f_{g_{eq}}=\sum_{i=(1-D_{e})/2}^{(D_{e}-1)/2}{p_{eq}(i)\times\delta(g_{eq}-\cos^{3}(\theta_{q}+\frac{2i}{LD_{e}}))},$ (99)
where
$p_{eq}(i)=Q(\frac{2i-1-LD\mu_{\varepsilon_{\theta_{q}}}}{LD\sigma_{\varepsilon_{\theta_{q}}}})-Q(\frac{2i+1-LD\mu_{\varepsilon_{\theta_{q}}}}{LD\sigma_{\varepsilon_{\theta_{q}}}}).$ (100)
According to (33), $\mathbb{G}_{a}=g_{ax}\times g_{ay}$ and applying (99), the PDF of $\mathbb{G}_{e}$ conditioned on $g_{et}$ can be determined as
$f_{\mathbb{G}_{e}|g_{et}}(\mathbb{G}_{e})=\sum_{i=(1-D_{e})/2}^{(D_{e}-1)/2}{\frac{p_{er}}{g_{et}}\times\delta(\frac{\mathbb{G}_{e}}{g_{et}}-\cos^{3}(\theta_{r}+\frac{2i}{LD_{e}}))}.$ (101)
The PDF of array factor $G_{e}$ is obtained by utilizing equations (99) and (101) as
[…]
$=\sum_{i=(1-D_{e})/2}^{(D_{e}-1)/2}\sum_{j=(1-D_{e})/2}^{(D_{e}-1)/2}{\int\frac{p_{et}(i)p_{er}(j)}{g_{et}}}$
×$\quad\times\delta(g_{et}-\cos^{3}(\theta_{t}+\frac{2i}{LD_{e}}))$
×$\quad\times\delta(\frac{\mathbb{G}_{e}}{g_{et}}-\cos^{3}(\theta_{r}+\frac{2j}{LD_{e}}))$
$=\sum_{i=(1-D_{e})/2}^{(D_{e}-1)/2}$ $\sum_{j=(1-D_{e})/2}^{(D_{e}-1)/2}{p_{et}(i)p_{er}(j)\delta(\mathbb{G}_{e}-q_{et}(i)q_{er}(j))},$
where
$q_{ek}(i))=\cos^{3}(\theta_{k}+\frac{2i}{LD_{e}}),\quad k\in\{t,r\}.$ (103)
We can rewrite result in (102) as
$f_{\mathbb{G}_{e}}(\mathbb{G}_{e})=\sum_{i=1}^{D_{e}^{2}}p_{e}(i)\delta(\mathbb{G}_{e}-q_{e}(i))$ (104)
where $\mathbf{q}_{e}=\mathbf{q}_{ex}\otimes\mathbf{q}_{ey}$ and $\mathbf{p}_{e}=\mathbf{p}_{ex}\otimes\mathbf{p}_{ey}$. Here, $\mathbf{q}_{ek}\in\mathbb{R}^{D_{e}\times 1}$ and $\mathbf{p}_{ek}\in\mathbb{R}^{D_{e}\times 1}$ where $k\in\{t,r\}$. $p_{ek}(i)$ and $q_{ek}(i)$ are given in (100) and (103), respectively.

### v3

@@ Appendix A Proof of Lemma 1
$\mu_{\varepsilon_{\theta_{q}}}=A_{\theta_{qx}}\mu_{x}+A_{\theta_{qy}}\mu_{y},$ (82a)
$\sigma^{2}_{\varepsilon_{\theta_{q}}}=A^{2}_{\theta_{qx}}\sigma^{2}_{x}+A^{2}_{\theta_{qy}}\sigma^{2}_{y}.$ (82b)
@@ Appendix B Proof of Lemma 3
The RVs $Z_{x}$ and $Z_{y}$, which are dependent on UMI fluctuations, follow a Gaussian distribution. From (36), the value of sector $i$, defined by $\frac{2i}{DN_{q}}<|Z_{x}|\leq\frac{2(i+1)}{DN_{q}}$, is $\frac{D^{2}(1-cos(\frac{2\pi i}{D}))}{2\pi^{2}i^{2}}$. The probability of $Z_{x}$ within this sector, $\text{Pr}(\frac{2i}{DN_{q}}<|Z_{x}|\leq\frac{2(i+1)}{DN_{q}})$ is calculated using the CDF of Gaussian-distributed $Z_{x}$, which can be expressed as $Q$-function. Consequently, the PDF of $g_{ax}$ will be given by
$f_{g_{ax}}=\sum_{i=0}^{lD-1}{p_{ax}(i)\times\delta(g_{ax}-\frac{D^{2}(1-cos(\frac{2\pi i}{D}))}{2\pi^{2}i^{2}})},$ (83)
@@
$p_{ax}(i)=Q(\frac{2i-DN_{q}\mu_{Z_{x}}}{DN_{q}\sigma_{Z_{x}}})$ $-Q(\frac{2(i+1)+DN_{q}\mu_{Z_{x}}}{DN_{q}\sigma_{Z_{x}}})$ (84)
$\qquad+Q(\frac{2i+DN_{q}\mu_{Z_{x}}}{DN_{q}\sigma_{Z_{x}}})$ $-Q(\frac{2(i+1)-DN_{q}\mu_{Z_{x}}}{DN_{q}\sigma_{Z_{x}}}).$
Note that the PDF of RV $g_{ay}$ can be derived similarly to (83) by replacing the subscript $x$ with $y$. According to (21), $\mathbb{G}_{a}=g_{ax}\times g_{ay}$ and applying (83), the PDF of array factor, $\mathbb{G}_{a}$ conditioned on the array factor in x-axis, $g_{ax}$ can be given as
@@
$\qquad\quad\sum_{i=0}^{lD-1}\frac{p_{ay}(i)}{g_{ax}}\times\delta(\frac{\mathbb{G}_{a}}{g_{ax}}-\frac{D^{2}(1-cos(\frac{2\pi i}{D}))}{2\pi^{2}i^{2}}),$
@@
$f_{\mathbb{G}_{a}}(\mathbb{G}_{a})$ $=\int f_{\mathbb{G}_{a}|g_{ax}}(\mathbb{G}_{a})f_{g_{ax}}dg_{ax}$ (86)
$=\sum_{i=0}^{lD-1}\sum_{j=0}^{lD_{a}-1}{\int\frac{p_{ax}(i)p_{ay}(j)}{g_{ax}}}$
$\quad\times\delta(\frac{\mathbb{G}_{a}}{g_{ax}}-\frac{D^{2}(1-\cos(\frac{2\pi i}{D}))}{2\pi^{2}i^{2}})$
$\quad\times{\delta(g_{ax}-\frac{D^{2}(1-\cos(\frac{2\pi j}{D}))}{2\pi^{2}j^{2}})}dg_{ax}$
$=\sum_{i=0}^{lD-1}\sum_{j=0}^{lD-1}{p_{ax}(i)p_{ay}(j)\delta(\mathbb{G}_{a}-q_{ax}(i)q_{ay}(j)))},$
@@
$q_{ak}(i)=\frac{D^{2}(1-\cos(\frac{2\pi i}{D}))}{2\pi^{2}i^{2}},\quad k\in\{x,y\}.$ (87)
@@
$f_{\mathbb{G}_{a}}(\mathbb{G}_{a})=\sum_{i=1}^{l^{2}D^{2}}p_{a}(i)\delta(\mathbb{G}_{a}-q_{a}(i)),$ (88)
where $\mathbf{q}_{a}=\mathbf{q}_{ax}\otimes\mathbf{q}_{ay}$ and $\mathbf{p}_{a}=\mathbf{p}_{ax}\otimes\mathbf{p}_{ay}$. Here, $\mathbf{q}_{ak}\in\mathbb{R}^{lD\times 1}$ and $\mathbf{p}_{ak}\in\mathbb{R}^{lD\times 1}$ where $k\in\{x,y\}$. $p_{ak}(i)$ and $q_{ak}(i)$ are given in (84) and (87), respectively.